# A Survey on Recent Advances in LLM-Based Multi-turn Dialogue Systems

### v1

@@ 2.2.2. UNiLM
UNILM (dong2019unified, ) is a unified pre-training model jointly optimized for multiple language modeling objectives with shared parameters, covering bidirectional, unidirectional, and sequence-to-sequence language models. The model undergoes pre-training using three types of language modeling tasks: unidirectional, bidirectional, and sequence-to-sequence prediction. Unified modeling is achieved through a shared Transformer network, incorporating specific self-attention masks to control the contextual conditions for predictions.
@@ 2.3. Encoder-decoder Transformer Architecture
The traditional Transformer model is built on an encoder-decoder architecture (raffel2020exploring, ), consisting of two Transformer blocks serving as the encoder and decoder, respectively. The encoder utilizes stacked multi-head self-attention layers to encode the input sequence and generate its latent representation. Meanwhile, the decoder performs cross-attention on these representations and autoregressively generates the target sequence. The schematic for this architecture is depicted in the left panel of Figure 1.
2.3.1. BART
@@
BART (lewis2019bart, ) is a denoising autoencoder designed for pre-training sequence-to-sequence models, combining Bidirectional and Auto-Regressive Transformers. It employs a sequence-to-sequence model applicable to diverse end tasks, utilizing a two-stage pre-training process involving text corruption and reconstruction. BART’s architecture, based on the standard Transformer for neural machine translation, can be seen as a generalization of BERT, GPT, and other recent pre-training schemes.
2.3.2. Text-to-Text Transfer Transformer
@@
T5 (raffel2020exploring, ), known as the ”Text-to-Text Transfer Transformer,” adopts the Transformer encoder-decoder structure, treating each text processing challenge as a ”text-to-text” task—generating new text from given input. This framework ensures consistent application of the same model, objective, training procedure, and decoding process across all tasks. The model employs an unsupervised span-corruption objective and a multi-task pre-training strategy. T5 establishes a unified model framework for various NLP tasks, simplifying the evaluation of different structures, pre-training objectives, and datasets across reading comprehension, summarization, and text classification tasks.
3. Fine-Tuning
3.1. Full Fine-Tuning
@@
Full Fine-Tuning (FFT) is a cornerstone technique in the domain of neural network adaptation, it involves optimizing all model parameters to integrate task-specific knowledge into the foundational architecture acquired during pre-training. FFT is instrumental in tailoring these models for specialized applications, ranging from nuanced language understanding to domain-specific tasks.
@@
FFT is based on optimizing a pre-trained model’s parameters. For a model $M$ characterized by parameters $\theta$, and a dataset $D$ designed for a specific task, the objective of FFT is to find the optimal parameter configuration $\theta^{*}$ that minimizes a defined loss function $\mathcal{L}$. This can be formally expressed as:
[…]
where $\mathcal{L}(M(\theta),D)$ represents the loss function that quantifies the disparity between the model’s predictions and the true outcomes in $D$.
@@
Selecting a pre-trained model that has learned from a wide variety of data begins the process. The next step involves preparing a dataset tailored to the specific task at hand. The core of FFT lies in thoroughly optimizing all model parameters to reduce the loss associated with the task. Additionally, the process includes enhancing the model’s learning through techniques such as data augmentation, advanced regularization, and optimizing the learning rate.
FFT is valued for its capacity to incorporate detailed, task-specific features into a model, thereby boosting its accuracy and effectiveness. This approach has contributed to significant advancements in fields like natural language processing, computer vision, and predictive analytics.
Despite FFT’s effectiveness in deeply adapting models, the development of neural network technologies has led to more focused fine-tuning methods that conserve resources. Parameter-Efficient Fine-Tuning (PEFT) methods, for example, adjust only a portion of the parameters, achieving a balance between thorough adaptation and computational demand, making model adaptation feasible even with limited resources.
3.2. Parameter-efficient Fine-Tuning
@@ 3.2.2. LoRA
3.2.3. Instruction Fine-Tuning
@@
3.3. Mitigating Fine-Tuning Instabilities
Fine-tuning pre-trained models, particularly in deep learning applications, often encounters instabilities leading to suboptimal performance(zhang2020revisiting, ; mosbach2021on, ). Instabilities manifest as erratic loss landscapes, difficulty in convergence, and sensitivity to hyperparameter settings. A common issue is representational collapse, where the model’s representations become less expressive over the course of fine-tuning (aghajanyan2020better, ). Researchers have developed various strategies to stabilize the fine-tuning process and enhance the robustness of fine-tuned models (dodge2020fine; mosbach2021on, ).
Aghajanyan et al. (aghajanyan2020better, ) present methods to mitigate representational collapse during fine-tuning of pre-trained models. They introduced robust representation through Regularization Fine-tuning (R3F) and Regularization and Reparameterization Fine-tuning (R4F), with the former adding regularization terms to the loss function and the latter extending R3F by introducing reparameterization. The approaches are formulated as:
@@ 3.3. Mitigating Fine-Tuning Instabilities
where $\mathcal{L}_{\text{original}}$ is the original loss function, $\lambda$ is the regularization strength, $\mathcal{R}$ is the regularization term, and Reparam is the reparameterization function.
Jiang et al. (jiang2019smart, ) propose the SMART (Smoothness-inducing Adversarial Regularization for Multitask Training) to enhance fine-tuning of pre-trained models through regularized optimization. It controls model complexity using Smoothness-inducing Adversarial Regularization and stabilizes updates with Bregman Proximal Point Optimization. The method is designed to improve robustness and efficiency in fine-tuning PLMs, addressing issues like overfitting and improving generalization in NLP tasks.
@@
where $\delta$ represents adversarial perturbations, $\rho$ balances the adversarial objective with regularization, and $f$ is the predictive function of the model.
Zhu et al. (zhu2019freelb, ) propose Free Large-Batch Adversarial Training(FreeLB), which is an adversarial training algorithm for improving the robustness and generalization of PLMs (zhu2019freelb, ). It enhances training by adding adversarial perturbations to word embeddings and minimizing adversarial risk.
@@
where $N$ is the batch size, $\delta_{i}$ are the adversarial perturbations, and $\rho$ balances the original loss with the adversarial term.
4. Prompt Engineering
Prompt engineering have gained significant attention in the field of helping PLMs understand a given task. Prompt engineering can be divided into two methods: Prompt Tuning and Tuning-free Prompting, which will be discussed in this section.
4.1. Prompt Tuning
Prompt tuning involves modifying the parameters of the pre-trained model or adjusting additional prompt-related parameters to enhance the adaptation of the pre-trained model to downstream tasks. A prominent approach in this regard is Pattern-Verbalizer-Pair (PVP) structure, which is initially proposed by (schick2020exploiting, ). Subsequent Prompt Tuning methods largely build upon the PVP framework. As shown in Figure 6, Gao et al. (gao2020making, ) has proved that the choice of templates and labels can lead to substantial differences in final accuracy. Therefore, current research primarily focuses on how to select or construct appropriate Patterns and Verbalizers. The design of prompts can be categorized into discrete prompts and continuous prompts.
Figure 6. The impact of templates and label words on prompt tuning.
4.1.1. Discrete Prompts
Previous efforts have been devoted to the exploration of discrete prompts, also referred to as hard prompts, which usually correspond to readable language phrases. The word embeddings of discrete template remain unchanged during training and won’t introduce any new parameter to the LLMs. In the following sections, we provide a detailed overview of several methods proposed for this purpose:
D1:Pattern-Exploiting Training
Pattern-Exploiting Training (PET) (schick-schutze-2021-exploiting, ) constructs a prompt set by manually generating PVP content p = (P, v), where $P$ is a function that takes $x$ as input and outputs a phrase or sentence $P(x)\in V^{*}$ that contains exactly one mask token. $v$ is an injective function $\mathcal{L}\rightarrow V$ that maps each label to a word in the masked language model vocabulary $M$. The PET can be formulated by:
(11) $s_{p}(l\ |\ x)=M(v(l)\ |\ P(x)),$
(12) $q_{p}(l\ |\ x)=\frac{e^{s_{p}(l|x)}}{{\textstyle\sum_{l^{\prime}\in\mathcal{L}}e^{s_{p}(l^{\prime}|x)}}},$
where $s_{p}$ is the score for label $l\in\mathcal{L}$, $q_{p}$ is the probability distribution over labels using softmax. Then, the cross-entropy between $q_{p}(l\ |\ x)$ and the true (one-hot) is used as the loss function for fine-tuning $M$ in p.
D2:LM-BFF
Finding the right prompts is cumbersome and fallible, which requires both domain expertise and understanding of NLP. Even if significant effort is invested, manual prompts are likely to be suboptimal. Therefore, LM-BFF (gao2020making, ) automatically generates prompts, including pruned brute-force searches to identify the best working label words, and utilizes the T5 model to automatically generate templates (raffel2020exploring, ). Given a fixed template $T(x)$, constructing a pruned set $V_{c}\subset V$ of the top $k$ vocabulary words based on their conditional likelihood using the initial $L$ can be formulated as:
(13) ${}_{v\in V}^{{Top-k}}\left\{\sum_{x_{in}\in D_{train}^{c}}\log{P_{\mathcal{L}}([MASK]=v\ |\ T(x_{in}))}\right\},$
where $P_{\mathcal{L}}$ denotes the output probability distribution of $\mathcal{L}$.
D3:Prompt Tuning with Rules
Prompt Tuning with Rules (PTR) (han2022ptr, ) can apply logical rules to automatically compose a small number of manually created sub-prompts into final task-specific prompts. Take relation classification as an example, given a sentence $x=\{...e_{s}...e_{o}...\}$, where $e_{s}$ and $e_{o}$ are the subject entity and the object entity respectively, the sub-prompt template and label word set of $e_{s}$ and $e_{o}$ can be formalized as:
(14) $T_{e_{s}/e_{o}}=``x\ the\ [MASK]\ e_{s}/e_{o}",$
$V_{e_{s}/e_{o}}=\{``person",``organization",...\},$
and the sub-prompt of relationship between $e_{s}$ and $e_{o}$ can be formalized as:
(15) $T_{e_{s}/e_{o}}=``x\ e_{s}\ [MASK]\ e_{o}",$
$V_{e_{s}/e_{o}}=\{``^{\prime}s\ parent\ was",``was\ born\ in",...\},$
by aggregating the sub-prompts, the complete prompt is as follows,
(16) $T=``x\ the\ [MASK]_{1}\ e_{s}\ [MASK]_{2}\ the\ [MASK]_{3}\ e_{o}",$
$V_{[MASK]_{1}}=\{``person",``organization",...\},$
$V_{[MASK]_{2}}=\{``^{\prime}s\ parent\ was,``was\ born\ in",...\},$
$V_{[MASK]_{3}}=\{``person",``organization",...\},$
the final learning objective of PTR is to maximize
(17) $\frac{1}{|X|}\sum_{x\in X}\log{\prod_{j=1}^{n}}p\left([MASK]_{j}=\phi_{j}(y)|T(x)\right),$
where $\phi_{j}(y)$ is to map the class $y$ to the set of label words $V_{[MASK]_{j}}$ for the $j$-th masked position $[MASK]_{j}$.
D4:Knowledgeable Prompt-tuning
Knowledgeable prompt-tuning (KPT) (hu2021knowledgeable, ) generate a set of label words for each label using external knowledge bases (KBs). The expanded label words are not simply synonyms of each other, but cover different granularities and perspectives, thus are more comprehensive and unbiased than the class name. Then, refinement methods are proposed to cope with the noise in the generated label words. This preserves high-quality words that can be used to fine tune PLMs and demonstrate the effectiveness of in-context learning (ICL).
4.1.2. Continuous Prompts
Instead of creating readable language phrases, continuous prompts convert the prompts into continuous vectors. Continuous prompts have their own parameters that can be tuned based on training data from downstream tasks.
C1:P-tuning
P-tuning (liu2023gpt, ) is an early attempt to achieve prompt tuning by employing continuous prompts. Instead of readable templates in discrete prompts, P-tuning uses continuous prompt embedding $p_{i}$ to build the prompt template:
(18) $T=\{[P_{0:i}],x,[p_{(i+q:j)}],y,[P_{j+1}:k]\},$
and then leverages an extra embedding function $f:[p_{i}]\to h_{i}$ to map the template to:
(19) $\{h_{0},...,h_{i},e(x),h_{(i+1)},...h_{j},e(y),h_{j+1},...j_{k}\},$
with the function $f$, task loss function can be optimized by updating the embedding $\{P_{i}\}_{i=1}^{k}$.
C2:Prefix Tuning
Prefix tuning(li2021prefix, ) freezes the parameters of PLMs and only optimizes the prefix parameters. Consequently, we only need to store the prefix for each task, making prefix-tuning modular and space-efficient. Given a trainable prefix matrix $M_{\phi}$ and a fixed pre-trained LM parameterized by $\theta$, the training objective is the same as that of full fine-tuning:
(20) $\max_{\phi}{\log{P_{\phi}(y|x)}=\max_{\phi}\sum_{i\in Y_{idx}}\log{P_{\phi}(z_{i}|h_{<i})}},$
where $h_{<i}$ is the concatenation of all neural network layers at time step $i$. If $i\in P_{idx}$, $h_{i}=P_{\theta}[i,:]$, otherwise, $h_{i}=LM_{\phi}(z_{i},h_{<i})$.
Lester et al. propose prompt tuning (lester2021power, ), which can be regarded as a simplification of prefix tuning. It greatly reduces the number of parameters and for the first time demonstrates that using only prompt tuning is also competitive.
Figure 7. From P-tuning to P-tuning v2. (liu2021p, )
C3: P-tuning V2
P-tuning v2 (liu2021p, ) is an implementation of prefix-tuning (li2021prefix, ) and mixture of soft prompts (qin2021learning, ). As shown in Figure 7, when compared to P-tuning, it not only has more tunable task-specific parameters (from 0.01% to 0.1%-3%) to allow more per-task capacity while being parameter-efficient, but also adds prompts to deeper layers have more direct impact on model predictions. P-tuning v2 is always comparable to fine-tuning at all scales but with only 0.1% task-specific parameters needed comparing to fine-tuning, proving that prompt-tuning can effectively assist PLMs in adapting to downstream tasks.
4.2. Tuning-free Prompting
@@
Tuning-free Prompting directly generate answers without modifying the parameters of the PLMs. These can optionally leverage response prompts to enhance inputs, previous studies have explored the impact of prompts on the generation effectiveness of PLMs and have provided numerous tricks for creating prompts. The mainstream tuning-free prompting methods include ICL and Chain-of-thought (CoT).
@@ 4.2.1. In-Context Learning
ICL uses multiple input-output demonstration pairs to PLMs to generate the desired response, which is first proposed along with GPT-3. ICL is an effective method as no parameter is tuned, an example prompt of ICL is shown in the Figure 8.
@@ 4.2.2. Chain-of-Thought
CoT (wei2022chain, ) improves performance on a range of arithmetic, commonsense, and symbolic reasoning tasks by mimics a step-by-step thought process for arriving at the answer. An example prompt of CoT is shown in the Figure 8. Zero-shot CoT (kojima2022large, ) is a classical CoT method that guides the model in reasoning and generating results by employing the same prompt ”Let’s think step by step,” across different tasks. Another classical approach is Least-to-Most (zhou2022least, ), which decomposes the target problem into a series of simpler sub-problems, solving them gradually.
@@ 5. LLM Based Task-oriented Dialogue Systems
The TOD system assists users in achieving specific domain-related goals, such as hotel reservations or restaurant queries, through interactive conversations. Due to its pronounced utility, this technology has garnered increasing attention from researchers in recent years. Generally speaking, TOD can be divided into pipeline-based TOD and end-to-end TOD. In this section, we will give a comprehensive introduction to LLM-based TOD.

### v2

@@ 2.2.2. UNiLM
UNiLM (dong2019unified, ) is a unified pre-training model optimized for multiple language modeling objectives, including bidirectional, unidirectional, and sequence-to-sequence tasks. It uses a shared Transformer network with specific self-attention masks to control contextual conditions for predictions, making it versatile for different language modeling tasks.
@@ 2.3. Encoder-decoder Transformer Architecture
The traditional Transformer model (raffel2020exploring, ) consists of an encoder-decoder architecture, where the encoder uses multi-head self-attention to process the input sequence and generate its latent representation, while the decoder autoregressively generates the target sequence through cross-attention.
@@
BART (lewis2019bart, ) is a denoising autoencoder that combines bidirectional and autoregressive Transformers. It employs a two-stage pre-training process involving text corruption and reconstruction, making it a versatile model for various sequence-to-sequence tasks. BART can be seen as a generalization of BERT, GPT, and other pre-training approaches.
@@
T5 (raffel2020exploring, ), or the ”Text-to-Text Transfer Transformer,” applies the encoder-decoder structure to treat all NLP tasks as text-to-text problems. Using a span-corruption objective and a multi-task pre-training strategy, T5 provides a unified framework that simplifies the application of the same model, training procedure, and decoding process across a wide range of tasks such as reading comprehension, summarization, and text classification.
2.4. Fine-Tuning
2.4.1. Full Fine-Tuning
@@
Full Fine-Tuning (FFT) is a key technique in neural network adaptation, involving the optimization of all model parameters to integrate task-specific knowledge into a pre-trained model. It is used to tailor models for specialized tasks, such as language understanding and domain-specific applications.
@@
FFT optimizes the parameters of a pre-trained model $M$ with parameters $\theta$ using a dataset $D$. The goal is to find the optimal parameters $\theta^{*}$ that minimize the loss function $\mathcal{L}$, as expressed by:
(1) $\theta^{*}=\underset{\theta}{\mathrm{argmin}}\ \mathcal{L}(M(\theta),D),$
where $\mathcal{L}(M(\theta),D)$ measures the model’s prediction error.
@@
FFT begins with selecting a pre-trained model, followed by preparing a task-specific dataset. The core of FFT is optimizing all model parameters to minimize task loss, incorporating techniques like data augmentation, regularization, and learning rate tuning.
FFT is valuable for incorporating task-specific features, significantly improving accuracy and performance. However, the emergence of Parameter-Efficient Fine-Tuning (PEFT) methods, which adjust only part of the model’s parameters, offers a more resource-efficient alternative while maintaining performance.
2.4.2. Parameter-efficient Fine-Tuning
@@
Instruction Fine-Tuning
@@
2.4.3. Mitigating Fine-Tuning Instabilities
Fine-tuning pre-trained models often encounters instabilities that hinder convergence and degrade performance (zhang2020revisiting, ; mosbach2021on, ). These challenges, such as erratic loss landscapes and representational collapse (aghajanyan2020better, ), have spurred research into methods for stabilizing fine-tuning processes and enhancing robustness (dodge2020fine, ; mosbach2021on, ).
Aghajanyan et al. (aghajanyan2020better, ) address representational collapse with Regularization Fine-tuning (R3F) and Regularization and Reparameterization Fine-tuning (R4F). R3F introduces a regularization term to the loss function, while R4F extends it by incorporating reparameterization. These are expressed as:
@@ 2.4.3. Mitigating Fine-Tuning Instabilities
where $\mathcal{L}_{\text{original}}$ is the task-specific loss, $\lambda$ controls regularization strength, $\mathcal{R}$ is the regularization term, and Reparam represents the reparameterization function.
Jiang et al. (jiang2019smart, ) propose SMART (Smoothness-inducing Adversarial Regularization for Multitask Training), which combines smoothness-inducing adversarial regularization with Bregman Proximal Point Optimization to stabilize fine-tuning and improve generalization. The method is formulated as:
@@
where $\delta$ are adversarial perturbations, $\rho$ balances regularization, and $f$ is the model’s predictive function.
Zhu et al. (zhu2019freelb, ) introduce FreeLB, an adversarial training approach that enhances robustness and generalization by injecting adversarial perturbations into word embeddings during training. It minimizes adversarial risk with:
@@
where $N$ is the batch size, $\delta_{i}$ are perturbations, and $\rho$ controls the trade-off between task loss and adversarial regularization.
These approaches collectively aim to stabilize the fine-tuning process, addressing challenges like overfitting and sensitivity to hyperparameters while improving model robustness and task-specific performance.
2.5. Prompt Engineering
@@
Tuning-free Prompting directly generate answers without modifying the parameters of the PLMs. These can optionally leverage response prompts to enhance inputs, previous studies have explored the impact of prompts on the generation effectiveness of PLMs and have provided numerous tricks for creating prompts. The mainstream tuning-free prompting methods include In-context Learning (ICL), Chain-of-thought (CoT) and Retrieval Augmented Generation (RAG).
@@ 2.5.1. In-Context Learning
ICL uses multiple input-output demonstration pairs to PLMs to generate the desired response, which is first proposed along with GPT-3.
@@ 2.5.2. Chain-of-Thought
CoT (wei2022chain, ) improves performance on a range of arithmetic, commonsense, and symbolic reasoning tasks by mimics a step-by-step thought process for arriving at the answer. Zero-shot CoT (kojima2022large, ) is a classical CoT method that guides the model in reasoning and generating results by employing the same prompt ”Let’s think step by step,” across different tasks. Another classical approach is Least-to-Most (zhou2022least, ), which decomposes the target problem into a series of simpler sub-problems, solving them gradually. Furthermore, the latest studies (coe, ) invesitigate to equip CoT with exemplar to provide structured guidance for better reasoning.
@@
2.5.3. Retrieval Augmented Generation
LLMs are facing the knowledge boundary problem, which stems from the fixed and static knowledge during training, making it difficult for LLMs to respond accurately to queries about recent events or niche domains. RAG addresses these issues by incorporating an external retrieval mechanism to enhance the generative process (lewis2020retrieval, ). By leveraging a retriever module, RAG accesses relevant documents from external knowledge bases, ensuring that the information used during generation is both up-to-date and contextually appropriate.
The development of RAG can be divided into three stages: Naive RAG, Advanced RAG and Modular RAG. Naive RAG (ma2023query, ) simply retrieves knowledge by calculating the similarity between the query and the database, and enhances the model’s generation based on the retrieved knowledge. Advanced RAG introduces pre-retrieval and post-retrieval strategies to overcome the limitations of Naive RAG. For example, Peng et al. (peng2024large, ) utilize LLMs to rewrite the long-tail queries. Modular RAG is composed of various modules, such as a retrieval module and a re-ranking module, allowing the Modular RAG pipeline to be conveniently restructured to adapt to different task requirements (yu2022generate, ; shao2023enhancing, ).
@@ 3. LLM Based Task-oriented Dialogue Systems
Table 2. Recent Advances in Task-Oriented Dialogue Systems.
Structure Task Methods Description
Pipeline-based NLU Zero-Shot-Adapters(comi2023zero, ) An intent detection method based on BERT model.
ICDA(lin2023selective, ) An approach based on LLMs and pointwise V-information.
Span-ConveRT(coope2020span, ) Integrates conversational knowledge coded in LLMs.
LEONA(siddique2021linguistically, ) Employs LLMs to provide contextualized representations.
COF-COT(nguyen2023cof, ) Breaks down NLU tasks into multiple reasoning steps.
DST SUMBT(lee2019sumbt, ) Learns the relations of slots through attention mechanisms.
IC-DST(hu2022context, ) A zero-shot and few-shot DST framework based on ICL.
SAVN(wang2020slot, ) Shares knowledge between slots and input to determine slots.
MinTL(lin2020mintl, ) Allows to plug-and-play pre-train seq2seq models.
LDST(feng2023towards, ) A DST framework leveraging fine-tuned LLaMa model.
RefPyDST (king2023diverse, ) Re-frames DST task as a Python programming task.
IDIC-DST (yi2024intent, ) Extract user’s intents to better retrieve examples.
FaS-DST (feng2024fact, ) Re-frames DST task as a dialogue summarization task.
PL GPT-Critic(jang2022gpt, ) Applies the fine-tuned GPT-2 model to address PL tasks.
CASPl(ramachandran2021causal, ) Infers human intentions rather than imitate demonstrations.
Galaxy(he2022galaxy, ) Introduces a new action prediction task during pre-training.
GDP-Zero(yu2023prompt, ) Prompts LLMs to act as a policy prior to solve PL tasks.
NLG SC-GPT(peng2020few, ) Pre-trained on a large set of annotated NLG corpus.
Few-Shot-NLG(chen2019few, ) Compose coherent sentences based on prior knowledge.
STs(baheti2020fluent, ) Identifies the best answer from candidate responses.
Controllable NLG(qian2022controllable, ) Utilizes prefix tuning method to address NLG tasks.
End-to-End Fully end-to-end Soloist (peng2021soloist, ) Divided into only three sub-tasks to reduce data overhead.
SimpleTOD(hosseini2020simple, ) A method based on LLMs trained on all sub-tasks.
UBAR (yang2021ubar, ) Fine-tunes the GPT-2 on the entire dialog session.
SyncTOD(saley-etal-2024-synergizing, ) Generates system responses in a single step.
Modularly end-to-end PPTOD(su2021multi, ) Pre-trained with four TOD tasks and enhanced by prompts.
SPACE series (he2022galaxy, ; he2022space, ; he2022unified, ) Pre-trained on limited labeled and abundant unlabeled data.
SynTOD(chieu-etal-2024-syntod, ) Generate synthesis responses to fine-tune LLMs.
The TOD system is designed to assist users in achieving specific goals within a particular domain, such as making hotel reservations or querying restaurant menus, through interactive conversations. Due to its practical applications, TOD has garnered significant attention from researchers, leading to the development of various methods.
TOD systems powered by LLMs excel at accurately recognizing user intent and efficiently managing task-specific details. They are particularly effective in multi-turn dialogue, maintaining coherence and ensuring appropriate responses throughout the conversation. Furthermore, LLMs’ ability to handle cross-domain tasks enables TOD systems to seamlessly switch between different activities, such as booking a flight or providing local recommendations, thereby enhancing the overall user experience.
TOD systems are generally classified into two main categories: pipeline-based and end-to-end systems. Pipeline-based systems break down the dialogue process into multiple stages, including natural language understanding, dialogue state tracking, and response generation. In contrast, end-to-end systems aim to map user inputs directly to responses in a unified model. A prominent real-world application of TOD is in virtual assistants, such as those used for booking services by platforms like Google Assistant or Amazon Alexa, where LLMs enable more fluid, context-aware interactions.
3.1. Pipeline-based Methods